# GraspClutter6D: A Large-scale Real-world Dataset for Robust Perception and Grasping in Cluttered Scenes

### v1

@@ V CONCLUSION
We presented GraspClutter6D, a dataset for robotic grasping and perception in real-world cluttered environments. Benchmark evaluations showed that existing methods face performance gaps in cluttered scenes, particularly when handling occlusion. Networks trained on our dataset achieved improved grasping performance, highlighting its value as a training resource. By publicly releasing our dataset and toolkits, we hope this dataset will serve as a foundation for advancing robotic manipulation in cluttered environments.
ACKNOWLEDGMENT
This work was supported by the Technology Innovation Program(00442029, Development of Tactile Intelligence in Robotic Hands Based on Tactile Data Learning to Manipulate Irregular Multiple Types of Objects) funded by the Ministry of Trade Industry & Energy(MOTIE, Korea)

### v3

@@ V CONCLUSION
We presented GraspClutter6D, a dataset for robotic grasping and perception in real-world cluttered environments. Our benchmark evaluations revealed a significant gap, as existing state-of-the-art methods show degraded performance in our high-complexity scenes. We demonstrated that GraspClutter6D serves as an effective training resource; models trained on our dataset achieved substantially improved grasping and perception, highlighting its value in developing robust systems to close this gap. For future work, we will focus on gripper-aware and reactive grasping motions, as we observed that simple approach trajectories often cause failures despite accurate grasp pose detection. By publicly releasing our dataset and toolkits, we hope this dataset will serve as a foundation for advancing robotic manipulation in cluttered environments.